Smithen Company, a wholesale distributor, has been operating for only a few months. The company sells three products—sinks, mirrors, and vanities. Budgeted sales by product and in total for the coming month are shown below based on planned unit sales as follows:

Units Percentage
  Sinks 600 50%
  Mirrors 300 25%
  Vanities 300 25%
  Total 1,200 100%
Product
Sinks Mirrors Vanities Total
  Percentage of total sales 50% 30 20 100%
  Sales $240,790.00 100% $145,200 100% $98,010.00 100% $484,000.00 100%
  Variable expenses 72,237.00 30% 116,160 80% 53,905.50 55% 242,302.50 50%
  Contribution margin $168,553.00 70% $29,040 20% $44,104.50 45% $241,697.50 50%
  Contribution margin per unit 280.92 96.80 147.02
  Fixed expenses 209,000.00
  Operating income 32,697.50

Break-even point in unit sales :

fixed expenses/overall CM ratio = 209,000/0.5 = 418,000

Break-even point in unit sales:

total fixed expenses/weighted-average CM per unit = 209,000/201.41* = 1,037.66 units

*(280.92 x 0.5) + (96.80 x 0.25) + (147.02 x 0.25)

Assume that actual sales for the month total $582,030 (1,400 units), with the CM ratio and per unit amounts the same as budgeted. Actual fixed expenses are the same as budgeted, $209,000. Actual sales by product are as follows: sinks, $196,645 (490 units); mirrors, $271,040 (560 units); and vanities, $114,345 (350 units).

1. Prepare a contribution format income statement for the month based on actual sales data. (Round your percentage answers to the nearest whole number.)

SMITHEN COMPANY

Contribution Margin Income Statement
Product
Sinks Mirrors Vanities Total
Percentage of total ? % ? % ? % ? %
Sales ? ? % ? ? % ? ? % ? ? %
variable expenses ? ? % ? ? % ? ? % ? ? %
contribution expenses $0 0 % $0 0 % $0 0 % 0 0 %
fixed expenses ?
Operating income (loss) $0

2. Compute the break-even point in sales dollars for the month, based on the actual data. (Round your intermediate calculations to the nearest whole percent. Round your final answer to the nearest whole dollar.)

3. Calculate the break-even point in unit sales for the month, based on the actual data. (Round your final answer to the nearest whole number.)
Answer #1

Let us compute the percentage of total sales.

  • Percentage = Actual sales/Total sales
  • Sinks =$196,645/$582,030 =34%
  • Mirrors = $271,040/$582,030 =46%
  • Vanities =$114,345/$582,030 =20%

​​​​​​​Let us compute the variable expenses.

  • Variable expenses= Sales × cost ratio
  • Sinks = $196,645 ×30%= $58,994.
  • Mirrors =$ 271,040 ×80%=$216,832.
  • Vanities=$114,345×55% =$62,890.

Let us compute the contribution margin

  • Contribution margin = Sales- variable expenses
  • Sinks =$196,645-$58,994 =$137,651.
  • Mirrors =$271,040-$216,832 =$ 54,208
  • Vanities= $114,345-$62,890=$51,455

​​​​​​

​​​​​Contribution margin income statement.

Sinks % Mirrors % Vanities % Total %
Percentage of total 34% 46% 20% 100%
Sales $196,645 100% $271,040 100% $114,345 100% $582,030 100%
Variable expenses $58,994 30% $216,832 80% $62,890 55% $338,716 58%
Contribution expenses $137,651 70% $54,208 20% $51,455 45% $243,314 42%
Contribution margin per unit $280.9 $96.8 $147
Fixed expenses $209,000
Operating income (loss) $34,314

2. Break-even point in sales dollars = Fixed costs / Weighted average contribution margin ratio

= $209,000/0.42

=$497,619.

3. Calculation of Break even point in unit sales.

weighted average contribution margin per unit= (Contribution margin per unit × sales mix)

Sinks = $208.9×34% =71.026

Mirrors =$96.8×46%=44.528

Vanities=$147×20% =$29.4

Total =$144.954 ($145).

Break even point in unit sales = Fixed cost/ weighted average contribution margin per unit

=$209,000/$145

=1,441 units.

____×____

All the best

Any doubt comment below

Please do support.

​​​​​​